A particle of mass m thrown vertically upward with velocity v0. The air resistance is mgcv² where c is a constant with v us the velocity at any time t. Show that the time taken by bthe particle to reach the highest point is given by v0 underroot c =tan(gt underroot c)


1
Expert's answer
2021-03-04T23:45:39-0500

Upward motion

Using Newton’s 2nd law gives:


"my''=-mgcy'^2-mg"

"\dfrac{dv}{dt}=-gcv^2-g"

"\int\dfrac{dv}{cv^2+1}=-\int g dt"

"\dfrac{1}{\sqrt{c}}\tan^{-1}(\sqrt{c}v)=-gt+C_1"

"t=0:C_1=\dfrac{1}{\sqrt{c}}\tan^{-1}(\sqrt{c}v_0)""\dfrac{1}{\sqrt{c}}\tan^{-1}(\sqrt{c}v)=-gt+\dfrac{1}{\sqrt{c}}\tan^{-1}(\sqrt{c}v_0)"

When the particle reaches the highest point "v=0"


"0=-gt_{vertex}+\dfrac{1}{\sqrt{c}}\tan^{-1}(\sqrt{c}v_0)"

"t_{vertex}=\dfrac{1}{g\sqrt{c}}\tan^{-1}(\sqrt{c}v_0)"